Sentry Manu. just paid a dividend of $5 per share. The dividend is expected to grow...

  1. Sentry Manu. just paid a dividend of $5 per share. The dividend is expected to grow at a constant rate of 6% per year. The price of Sentry Manu. stock today is $32 per share. If Sentry Manu. decides to issue new common stock, flotation costs will equal $2.30 per share. Sentry Manu. marginal tax rate is 35% above information, the cost of new common stock (external equity) is

    1. 23.19%

    2. 25.73%

    3. 23.85%

    4. 24.23%
Answer #1

For new common stock, price needs to be adjusted for flotation cost,

Price adjusted for flotation cost = $32 - $2.30 = $29.70

Div1 = $5 * (1 + 6%) = $5.30

r - 0.06 = 5.30/29.70

r = 0.1785 + 0.06

r = 23.85%